OpenStudy (anonymous):

If angle ABC = angle CBD, find x. 5 -10 10 2/5 I got 10 but do u think its -10 ?

Directrix (directrix):

Diagram?

Directrix (directrix):

Okay. Let me crank it out and see.

Directrix (directrix):

@Addicto_ Do you remember what equation you used to get the 10?

OpenStudy (anonymous):

4x + 30 = 5x + 20 5x - 4x = 30 - 20 x = 10

OpenStudy (anonymous):

this is how i got it

Directrix (directrix):

Yes, x = 10 but not -10 as your work shows.

OpenStudy (anonymous):

ok :)

Directrix (directrix):

Alrighty.